The perimeter of a pasture is 1250 feet. The pasture is four times longer than it is wide. Find the dimensions by setting up a system of equations.
Answer by askmemath(368)

Let width = W
then length = 4w
Perimeter = 2(l+b) = 2(w+4w) = 2(5w) = 10w
But we know this = 1250
So 10w = 1250
w = 125
L = 4x125 = 500